If K<1, ∆G
   <0
   =1
   >0
   =0

Calculate ∆Gº at 25ºC for the reaction where ∆Hº=23.7 kJ kJ and ∆Sº=52.4 J/K:
   -1290 kJ
   22.4 kJ
   -1.56x104kJ
   8.08 kJ
Answer #1

1)

∆G = -RT*ln K

K<1

So,

ln K < 0

So,

RT*ln K < 0

This means,

-RT*ln K > 0 and hence ∆G > 0

Answer: > 0

2)

ΔHo = 23.7 KJ

ΔSo = 52.4 J/K

= 0.0524 KJ/K

T= 25.0 oC

= (25.0+273) K

= 298 K

use:

ΔGo = ΔHo - T*ΔSo

ΔGo = 23.7 - 298.0 * 0.0524

ΔGo = 8.08 KJ

Answer: 8.08 KJ